# An Information Density-based Machine-Generated Text Detector

### v2

@@ Abstract
The Uniform Information Density (UID) principle posits that humans prefer to spread information evenly during language production. We examine if this UID principle can help capture differences between Large Language Models (LLMs)-generated and human-generated texts. We propose GPT-who, the first psycholinguistically-aware multi-class domain-agnostic statistical detector. This detector employs UID-based features to model the unique statistical signature of each LLM and human author for accurate authorship attribution. We evaluate our method using 4 large-scale benchmark datasets and find that GPT-who outperforms state-of-the-art detectors (both statistical- & non-statistical) such as GLTR, GPTZero, DetectGPT, OpenAI detector, and ZeroGPT by over $20$% across domains. In addition to better performance, it is computationally inexpensive and utilizes an interpretable representation of text articles. We find that GPT-who can distinguish texts generated by very sophisticated LLMs, even when the overlying text is indiscernible. UID-based measures for all datasets and code are available at https://github.com/saranya-venkatraman/gpt-who.
@@
Saranya Venkatraman Penn State University saranyav@psu.edu                        Adaku Uchendu MIT Lincoln Laboratory adaku.uchendu@ll.mit.edu                        Dongwon Lee Penn State University dongwon@psu.edu
@@ 1 Introduction
To that end, we propose GPT-who, the first psycholinguistically-aware supervised domain-agnostic task-independent multi-class statistical-based detector. GPT-who calculates interpretable Uniform Information Density (UID) based features from the statistical distribution of a piece of text and automatically learns the threshold (using Logistic Regression) between different authors.
To showcase the detection capabilities of GPT-who, we use 4 large LLM benchmark datasets: TuringBench Uchendu et al. (2021), GPABenchmark Liu et al. (2023b), ArguGPT Liu et al. (2023a), and Deepfake Text in-the-wild Li et al. (2023). We find that GPT-who outperforms state-of-the-art statistical detectors and is at par with task and domain-specific fine-tuned LMs for authorship attribution. This performative gain is consistent across benchmark datasets, types of LLMs, writing tasks, and domains.
It is even more remarkable that this performative gain is accompanied by two essential factors: First, GPT-who is computationally inexpensive as it eliminates the need for any LLM fine-tuning. It utilizes a freely available off-the-shelf LM to compute token probabilities, followed by logistic regression using a small set of carefully crafted and theoretically motivated UID features. Second, GPT-who provides a means to interpret and understand its prediction behaviors due to the rich feature space it learns from. UID-based features enable observable distinctions in the surprisal patterns of texts, which help in understanding GPT-who’s decision-making on authorship (Figure 1).
We also analyze the UID distributions of different LLMs and human-generated texts across all datasets and find that humans distribute information more unevenly and diversely than models. In addition, UID features are reflective of differences in LLM architectures or families such that models that share architectures have similar UID distributions within but not outside their category. We find that UID-based features are a consistent predictor of authorship. Even when there are not glaring differences between uniform and non-uniform text, the differences in UID distributions are easily detectable and a powerful predictor of authorship, since they successfully capture patterns that go beyond the lexical, semantic, or syntactic properties of text. Our work indicates that psycholinguistically-inspired tools can hold their ground in the age of LLMs and a simpler theoretically-motivated approach can outperform complex and expensive uninterpretable black-box approaches for machine text detection.
@@ 2.1 Uniform Information Density (UID)
UID in human language production has been studied by measuring the amount of information content per linguistic unit (sentence length/number of words) or by studying any sudden changes in surprisal at the onset of a word or sentential element Xu and Reitter (2016); Jaeger and Levy (2007). A rich body of work in psycholinguistics has led to the finding that, in language production, humans try to spread information content or surprisal evenly and maintain UID through their lexical, syntactic, phonological, and semantic choices Frank and Jaeger (2008); Xu and Reitter (2018); Jaeger (2010); Mahowald et al. (2013); Tily and Piantadosi (2009).
@@ 2.2 Machine-Generated Text Detection
More recently, due to the increased ubiquity of LLMs, we need more interpretable, and less deep learning-based models. Deep learning models have been shown to be the most susceptible to adversarial perturbations than others Pu et al. (2022). To that end, we propose the first supervised statistical-based technique, that calculates UID-based features of a given text and uses a classical machine learning model to automatically decide thresholds.
@@ 3.1 UID-based features
    Mean Surprisal ($\mu$) of an article (y) defined as follow:
@@
Random No Spans Min + Max UID spans
Span Length (N) UID spans N=4 N=10 N=15 N=20 N=30
@@
Table 1: Max. & Min. UID spans ablation study: Setting a span length of N=20 tokens maximized performance across large-scale datasets (N>30 leads to subsequently lower and eventually consistent performance). It can be seen that our min/max features tremendously impact performance against randomly sampled or no span features at all.
@@ 4 Empirical Validation
We use Meister et al. (2021)’s implementation of UID-based scores [^1] https://github.com/rycolab/revisiting-uid/tree/main and use the publicly available off-the-shelf pre-trained GPT2-XL language model [^2] https://huggingface.co/gpt2-xl to obtain conditional probabilities. For all our experiments, we calculate the UID features for the publically released train and test splits of all datasets. We train a logistic regression model [^3] https://scikit-learn.org/stable/ using these features on the train splits and report performance on the test splits. We replicate all the original evaluation settings and metrics for each of the datasets (except one setting from the ArguGPT Liu et al. (2023a) dataset that required access to unreleased human evaluation data). We do this to be able to directly compare the performance of GPT-who with current state-of-the-art detection methods reported so far.
@@ 4.1 Datasets
To test the applicability of GPT-who across text detection tasks, we run all experiments across 4 large-scale and very recent datasets that span over 15 domains and 35 recent LMs.
@@ 4.3 UID Signatures of Authors
Given that humans tend to optimize UID, we study if different models spread surprisal in ways that are distinguishable from each other and human-written text and if we can observe unique UID signatures of different LM families. To this end, we plot the UID score distributions of different text generators across (see Figures 4, 4(a), and 4(b)). We observe that, generally, the UID scores of human-written text have a higher mean and larger standard deviation than most machine-written text across writing task types, domains, and datasets. This implies that human-written text tends to be more non-uniform and diverse in comparison to machine-generated text. Hence, machines seem to be spreading information more evenly or smoothly than humans who are more likely to have fluctuations in their surprisal distributions. Going a step further, if we compare models to other models, we see that models that belong to the same LM family by architecture tend to follow similar UID distribution. For example, in Figure 4, the dotted lines separate LMs by their architecture type and it can be seen, for example, that all GPT-2 based models have similar UID distributions, all Grover-based models have similarities, but these groups are distinct from each other. This indicates that UID-based features can capture differences in text generated by not only humans and models but also one step further to capture differences between individual and multiple models and LM families. To our knowledge, this is the first large-scale UID-based analysis of recent machine and human-generated text across writing tasks and domains.
4.4 Machine Text Detection Performance
Overall, GPT-who outperforms other statistical-based detectors and is at par with transformers-based fine-tuned methods for 2 out of 4 benchmarks. For GPABenchmark (Table 2), across all
@@ 4.4 Machine Text Detection Performance
task types and domains, GPT-who outperforms GPTZero, ZeroGPT, DetectGPT and, OpenAI’s detector by over 40%. The machine-generated texts for this task are from 7 very recent and highly sophisticated LLMs (including GPT3.5, GPT3 variants), making the detection of machine-generated text a much more challenging task on which GPT-who outperforms other detectors.
@@ 5 Discussion
We turn to the UID principle, which states that humans prefer to spread information evenly in language, to automatically extract features that measure the spread and flow of information content or surprisal in texts. Our UID-based features are formulated to capture how surprisal is distributed in an article as they measure the local and global variance, mean, and most uniform and non-uniform segments of a text. This rich and succinct representation space drives the predictive capability of our proposed detector and the interpretability of its representations. Analysis of this feature space reveals that human-written text tends to be more non-uniform in comparison to machine-generated text. Hence, machines seem to be spreading information more evenly or smoothly than humans who are more likely to have fluctuations in their surprisal distributions. We also find that UID-based features can capture differences between text generated by not only humans and models but also capture differences between multiple models and LM families. Our main contribution is a novel psycholinguistically-aware domain-agnostic multi-class statistical-based machine-generated text detector, GPT-who, that:
@@
  • Intuitively interpretable due to its psycholinguistically motivated UID-based feature space.
@@
Acknowledgments

### v3

@@ Abstract
The Uniform Information Density (UID) principle posits that humans prefer to spread information evenly during language production. We examine if this UID principle can help capture differences between Large Language Models (LLMs)-generated and human-generated texts. We propose GPT-who, the first psycholinguistically-inspired domain-agnostic statistical detector. This detector employs UID-based features to model the unique statistical signature of each LLM and human author for accurate detection. We evaluate our method using 4 large-scale benchmark datasets and find that GPT-who outperforms state-of-the-art detectors (both statistical- & non-statistical) such as GLTR, GPTZero, DetectGPT, OpenAI detector, and ZeroGPT by over $20$% across domains. In addition to better performance, it is computationally inexpensive and utilizes an interpretable representation of text articles. We find that GPT-who can distinguish texts generated by very sophisticated LLMs, even when the overlying text is indiscernible. UID-based measures for all datasets and code are available at https://github.com/saranya-venkatraman/gpt-who.
@@
Saranya Venkatraman The Pennsylvania State University saranyav@psu.edu                        Adaku Uchendu MIT Lincoln Laboratory adaku.uchendu@ll.mit.edu                        Dongwon Lee The Pennsylvania State University dongwon@psu.edu
@@ 1 Introduction
To that end, we propose GPT-who, the first psycholinguistically-inspired supervised domain-agnostic task-independent multi-class statistical-based detector. GPT-who calculates interpretable Uniform Information Density (UID) based features from the statistical distribution of a piece of text and automatically learns the threshold (using Logistic Regression) between different authors. Such features are calculated using the surprisals of tokens in a text, for example, the variance of surprisals, the mean, and so on (elaborated in Section 3.1). To showcase the detection capabilities of GPT-who, we use 4 large LLM benchmark datasets: TuringBench Uchendu et al. (2021), GPABenchmark Liu et al. (2023b), ArguGPT Liu et al. (2023a), and Deepfake Text in-the-wild Li et al. (2023). We find that GPT-who outperforms state-of-the-art statistical detectors and is at par with task and domain-specific fine-tuned LMs. This performative gain is consistent across benchmark datasets, types of LLMs, writing tasks, and domains.
In addition to improved detection performance, GPT-who is computationally inexpensive as it eliminates the need for any LLM fine-tuning. It utilizes a freely available off-the-shelf LM to compute token probabilities, followed by logistic regression using a small set of carefully crafted and theoretically motivated UID features. GPT-who also provides a means to interpret and understand its prediction behaviors due to the rich feature space it learns from. UID-based features enable observable distinctions in the surprisal patterns of texts, which help in understanding GPT-who’s decision-making on authorship (Figure 1).
We also analyze the UID feature distributions of different LLMs and human-generated texts for the 4 datasets and find that humans distribute information more unevenly and diversely than models. In addition, UID features are reflective of differences in LLM architectures or families such that models that share architectures have similar UID distributions within but not outside their category. We find that UID-based features are a consistent predictor of authorship. Even when there are no glaring differences between uniform and non-uniform text, the differences in UID distributions are easily detectable and a powerful predictor of authorship, since they capture patterns that go beyond the lexical, semantic, or syntactic properties of text. Our work indicates that psycholinguistically-inspired tools can hold their ground in the age of LLMs and a simpler theoretically-motivated approach can outperform complex and expensive uninterpretable black-box approaches for machine text detection.
@@ 2.1 Uniform Information Density (UID)
UID in human language production has been studied by measuring the amount of information content per linguistic unit (number of words) or by studying any sudden changes in surprisal at the onset of a word or sentential element Xu and Reitter (2016); Jaeger and Levy (2007). A rich body of work in psycholinguistics has led to the finding that, in language production, humans try to spread information content or surprisal evenly and maintain UID through their lexical, syntactic, phonological, and semantic choices Frank and Jaeger (2008); Xu and Reitter (2018); Jaeger (2010); Mahowald et al. (2013); Tily and Piantadosi (2009).
@@ 2.2 Machine-Generated Text Detection
More recently, due to the increased ubiquity of LLMs, we need more interpretable, and less deep learning-based models. Deep learning models have been shown to be the most susceptible to adversarial perturbations than others Pu et al. (2022). To that end, we propose a supervised statistical-based technique, that calculates UID-based features of a given text and uses a classical machine learning model to automatically decide thresholds.
@@ 3.1 UID-based features
    Mean Surprisal ($\mu$) of an article (y) of length |y| in number of tokens is defined as follows:
@@
Random No Spans Span Length (N) of Min/Max UID spans
Human v. UID spans N=4 N=10 N=15 N=20 N=30
@@
Table 1: Max. & Min. UID spans ablation study: Setting a span length of N=20 tokens maximized performance (F1 score) across large-scale datasets (N>30 leads to subsequently lower and eventually consistent performance). It can be seen that our min/max features tremendously impact performance against randomly sampled or no span features at all.
@@ 4 Empirical Validation
We use Meister et al. (2021)’s implementation of UID-based scores [^1] https://github.com/rycolab/revisiting-uid/tree/main and use the publicly available off-the-shelf pre-trained GPT2-XL language model [^2] https://huggingface.co/gpt2-xl to obtain conditional probabilities. For all our experiments, we calculate the UID features for the publically released train and test splits of all datasets of each of the 4 benchmarks as they were released by the dataset developers. We train a logistic regression model [^3] https://scikit-learn.org/stable/ using these features on the train splits and report performance on the test splits. We averaged performance over 3 different random seeds and set the number of maximum iterations hyperparameter to 10k after testing a set of values. We replicate all the original evaluation settings and metrics for each of the datasets (except one setting from the ArguGPT Liu et al. (2023a) dataset that required access to unreleased human evaluation data). We do this to be able to directly compare the performance of GPT-who with current state-of-the-art detection methods reported so far.
@@ 4.1 Datasets
To test the applicability of GPT-who across text detection tasks, we run all experiments across 4 large-scale and recent datasets that span over 15 domains and 35 recent LMs.
@@ 4.3 UID Signatures of Authors
Given that humans tend to optimize UID, we study if different models spread surprisal in ways that are distinguishable from each other and human-written text and if we can observe unique UID signatures of different LM families. To this end, we plot the UID score distributions of different text generators across (see Figures 4, 5(a), and 5(b)). We observe that, generally, the UID scores of human-written text have a higher mean and larger standard deviation than most machine-written text across writing task types, domains, and datasets. This implies that human-written text tends to be more non-uniform and diverse in comparison to machine-generated text. Hence, machines seem to be spreading information more evenly or smoothly than humans who are more likely to have fluctuations in their surprisal distributions. Going a step further, if we compare models to other models, we see that models that belong to the same LM family by architecture tend to follow similar UID distribution. For example, in Figure 4, the dotted lines separate LMs by their architecture type and it can be seen, for example, that all GPT-2 based models have similar UID distributions, all Grover-based models have similarities, but these groups are distinct from each other. This indicates that UID-based features can capture differences in text generated by different LM families. To our knowledge, this is the first large-scale UID-based analysis of recent machine and human-generated text across writing tasks and domains.
@@
4.4 Machine Text Detection Performance
Overall, GPT-who outperforms other statistical-based detectors and is at par with transformers-based fine-tuned methods for 2 out of 4 benchmarks. For GPABenchmark (Table 2), across all task types and domains, GPT-who outperforms GPTZero, ZeroGPT, DetectGPT and, OpenAI’s detector by over 40%. The machine-generated texts for this task are from 7 very recent and highly sophisticated LLMs (including GPT3.5, GPT3 variants), making the detection of machine-generated text a much more challenging task on which GPT-who outperforms other detectors.
@@ 4.4 Machine Text Detection Performance
GPT-who is a statistical-based approach that outperforms other statistical-based approaches but is unsurprisingly outperformed by fine-tuned methods in 2 out of 4 benchmarks. In the case of statistical-based methods, it is typically very hard to come close to fine-tuned performance as such methods rely only on derived properties of the text and do not utilize the full raw text in training as is the case in the latter Jawahar et al. (2020). Despite this, GPT-who can exceed fine-tuned LM performance by 10% for 2 benchmarks.
4.5 Running Time
We measured the time taken for the one-time training or fine-tuning and inference for 6 testbeds from the InTheWild Dataset (the largest of all our benchmarks). We compare the average running times of DetectGPT, BERT, GPT-who and a stylometric detector in Table 6 and find that GPT-who is the fastest as it eliminates the need for any LM fine-tuning and makes a single inference call per text sample. Other methods require either LM fine-tuning or multiple inference calls (for example, DetectGPT). This computational load is greater than a single forward inference pass through one LM (GPT2) followed by logistic regression which is what GPT-who requires.
Method One-Time Training Inference
DetectGPT >10 hours 60 sec
BERT $\sim$1.5 hours 2 sec
Stylometry $\sim$1.5 hours 2 sec
GPT-who 20 min 0.8 sec
Table 6: Average Running time over 6 testbeds from the InTheWild dataset.
@@ 5 Discussion
We turn to the UID principle, which states that humans prefer to spread information evenly in language, to automatically extract features that measure the spread and flow of information content or surprisal in texts. Our UID-based features are formulated to capture how surprisal is distributed in an article as they measure the local and global variance, mean, and most uniform and non-uniform segments of a text. This rich and succinct representation space drives the predictive capability of our proposed detector and the interpretability of its representations. Analysis of this feature space reveals that human-written text tends to be more non-uniform in comparison to machine-generated text. Hence, machines seem to be spreading information more evenly or smoothly than humans who are more likely to have fluctuations in their surprisal distributions. However, this finding does not imply that humans are not producing uniform text. It is important to note that our work cannot provide support for or refute the UID hypothesis which comes from psycholinguistic studies such as those in Section 2.1. Our work shows that, given our operationalization of UID based on prior works, machine text is relatively more uniform than human-written text. While this might seem contradictory to UID theory, it does not still disprove that humans are uniform in their language production.
We conjecture that this unexpected finding is because we use GPT-2’s probability distribution to calculate surprisal, which is potentially a poor approximation of the “human” probability distribution. A closer-to-human probability distribution might (or might not) show humans to be more uniform than machines, though this determination is not within the scope of this work. It is crucial to note that uniformity is relative, and while machines are more uniform under this operationalization, it would still be true that human text is uniform as per the human’s probability distribution (that we do not have access to and can only approximate using some LM distribution, for example, GPT2-XL in our case). UID theory does not make any predictions on where machine-generated text might lie in the uniform to non-uniform spectrum but only indicates that humans are arranging utterances evenly as per their own language distribution.
Irrespective of its alignment with what theory suggests, we find that the UID-based features are very useful in distinguishing authors, which is the focus of this work. This is an important consideration and helps disentangle the utility of UID-inspired features from the cognitive plausibility of those feature calculations or UID approximations.
Thus, this operationalization of UID does not imply that humans are “less human” or machines are “more human” as it is an approximation of a theory that states that humans are uniform as per their language distribution. It does not have any further implications for machine-generated text and is unable to predict what happens in the case of machines.
We find that UID-based features can capture differences between text generated by not only humans and models but also capture differences between multiple models and LM families. Our main contribution is a psycholinguistically-inspired domain-agnostic statistical-based machine-generated text detector, GPT-who, that:
@@
  • Interpretable due to its psycholinguistically motivated UID-based feature space.
@@
6 Conclusion
We propose GPT-who, a statistical-based machine-generated text detector that utilizes features inspired by the Uniform Information Density (UID) principle rooted in the observation that humans tend to evenly distribute information or surprisal in language. We leverage UID-based features, including variance of surprisals and minimally/maximally uniform sub-sequences extracted from the surprisal sequence generated by an off-the-shelf LM. We demonstrate that these features are highly effective in discerning machine-generated text from human-generated text as they capture nuances in how models and humans distribute surprisal in their texts. Our findings have implications for enhanced text authenticity assessment.